OpenStudy (anonymous):

does anybody knows the three ordered pairs of y=-6?

Directrix (directrix):

@marili Do you want to find three ordered pairs of the line y = -6?

OpenStudy (anonymous):

no of the equation

Directrix (directrix):

Wel, y = -6 is the equation of a line so I think we're talking the same thing.

OpenStudy (anonymous):

But do you know the 3 ordered pair? Im so lost

Directrix (directrix):

If we were graphing or finding points of the equation y = x + 6, then we would choose an x, say 2, and substitute that in the equation. y = x + 6 y = 2 + 6 y = 8 So, one ordered pair of the equation, a point on the linear graph of the equation would be (2,8). Do you agree? @marili

OpenStudy (anonymous):

hmm?? I still don't understand.

Directrix (directrix):

What is the procedure you use to find ordered pairs? We'll use it. @marili

OpenStudy (anonymous):

well i use the y=mx+b, it iis easier for me

Directrix (directrix):

Here's the equation of the line: y=-6. Now, @marili, what have you been doing next in working these problems?

OpenStudy (anonymous):

I do not even know how to set up the problem?

Directrix (directrix):

If you want to use this --> y=mx+b, we will have to know what m and b are. So, how do you write y = -6 in y = mx+ b form? @marili

OpenStudy (anonymous):

Im not sure but i think it will be y=mx-6

Directrix (directrix):

Okay. And, because there is no x term in the equation y = -6, then what you just posted would be y = 0*x -6. So whatever is the value of x is multiplied by 0 which gives zero. The bottom line is that it makes no difference what the value of x is. That is because y is -6 all the time. I'll do one point and you the next. I select x = 1. So y = 0*x -6 becomes y = 0*1 - 6 = 0 - 6 = -6. The ordered pair is ( 1, -6) * Your turn. Find a point on the line, any point will do if it is on the line. @marili

OpenStudy (anonymous):

ok i think i kind of get it now